# Beyond Fidelity: Explaining Vulnerability Localization of Learning-based Detectors

### 05 Jan 2024 v1

@@ 5.4.4. FLC Among Sequence-based Explainers
Fig 10 illustrates the FLC results of sequence-based explainers, it reveals that:
The FLC scores achieved by sequence-based explainers fall short of expectations. On the Fan dataset, when explaining TokenLSTM, Shap attains the best FLC score, while it only reaches 30% of Expectation as shown in Fig 10 (b.1). On the SARD dataset, Shap performs the best in FLC among the five explainers, however, the observed outcome diverges significantly from Expectation as depicted own in Fig 10 (a.1),(a.2),(a.3).
There is minimal correspondence between the fidelity scores and FLC. One example is that Shap continues to outperform other explainers in terms of FLC on the SARD dataset. However, it does not achieve the highest fidelity scores across all detectors as shown in Fig 7 (a.1),(b.1),(c.1).
Considering the aforementioned observations, it is uncertain whether explanation methods can reliably provide valuable insights. We hypothesize that DL-based detectors aim to learn the distinctions between vulnerable and normal samples, attempting to discern whether the content represents a vulnerability or normal code fragments. Given that VTSs exist in both vulnerable and non-vulnerable programs, DL-based detectors might not prioritize the importance of VTS in determining vulnerabilities. Hence, theoretically, explanation methods should achieve high FLC. As they focus on the differences between vulnerable and normal samples. However, experimental outcomes do not consistently align with this expectation. Furthermore, there appears to be no direct correlation observed among fidelity, TLC, and FLC values. As suggested by T Ganz (ACMExplainDiscovery, ), this discrepancy might stem from DL-based detectors being trained not to recognize vulnerabilities but artifacts encompassing not only VFS. These findings underscore the need for further analysis of DL-based detectors and explanation approaches, which will be discussed in Section 5.5.
@@ 5.5.1.1   Experimental Setting
Dataset and detector consideration. When sampling the dataset, we chose to sample from the SARD portion and not from the Fan dataset. This decision was influenced by concerns regarding the potential presence of vulnerability-unrelated code changes in patch commits, as discussed in (dataQuality, ). Such changes could result in inaccurate labeling of vulnerabilities within the dataset. Simultaneously, we solely employ function-level detectors for experimentation and refrain from using slice-level detectors. This decision is based on the fact that DeepWuKong, SySeVR, and VulDeePecker detectors conduct slicing and symbolization operation. This process potentially exacerbates the differences unrelated to vulnerabilities between vulnerable and normal samples. For example, during symbolization, variable names such as “data” or “databuffer” in the original samples may be symbolized as “VAR1” and “VAR2”. However, in different samples or between a vulnerability sample and its corresponding normal sample, these variables may be symbolized as different values, such as “VAR1” in the vulnerability sample and “VAR2” in the normal sample. Subsequent experiments (section 5.5.2) reveal that these textual disparities interfere to some extent with DL-based detectors.
@@
Dataset and detector selection. After sampling, the remaining vulnerable sample sizes in the sub-test set are shown in Table 3, indicating a considerable reduction in the dataset sizes for each type of vulnerability. The most notable reductions are observed in CWE-190 and CWE-400. On the sub-dataset, we first retrain the models for Reveal, IVDetect, Devign, and TokenLSTM. The comparison of the performance on the sub-test dataset between the previous and retrained models for each detector is depicted in Fig 11. We can observe that Reveal and IVDetect exhibit minor performance differences between the previous and retrained models across the six types of sub-test datasets. However, Devign shows performance declines in CWE-20, CWE-119, and CWE-787, while TokenLSTM specifically demonstrates reduced performance in CWE-125, CWE-190, and CWE-400. We speculate that the performance declines may be related to the quantity and quality of the training samples. Considering the requirement for similar performance between the previous and retrained models, the number of test samples, and the need to involve both graph-based and sequence-based methods, we opt to continue the experiments on portions of CWE-20 and CWE-119 samples for Reveal, IVDetect, and TokenLSTM.
@@
Table 3. Number of vulnerable functions of each dataset after filtering.
@@ 5.5.1.2   Experimental Results
The explanation results for Reveal, IVDetect, and TokenLSTM are depicted in Fig 1213 and  14 respectively. From the results, we can observe:
@@
DL-based detectors focus on differentiating vulnerable and normal samples, a characteristic also mirrored by gradient-, decomposition-, and surrogate-based explainers. Within Reveal, all model-related explainers demonstrate an ability to achieve higher FLC scores while concurrently registering decreased TLC. Regarding IVDetect, while the FLC score of DeepLift remains relatively modest without notable improvements, both GradCAM and GNN-LRP consistently maintain notably high FLC scores, exhibiting a slight increase, particularly evident in the retrained model. In the case of TokenLSTM, it is notable that all deployed explainers, with the exception of DeepLift, attain higher FLC scores in the retrained models. The aforementioned observations suggest that the retrained model indeed tends to place more emphasis on the VFS during predictions compared to the previous model. Additionally, explanation methods, beyond perturbation-based approaches, to some extent, can capture this particular aspect of the model. The results depicted in Table 6 reveal that in both the retrained and previous models, the significance of VFS surpasses that of VTS. Nevertheless, when examining their explanations in both retrained and previous models, their results may still not be considered ideal. This implies that they are still prone to selecting features unrelated to vulnerabilities, which could be associated with both the deep learning models and the explanation methods themselves. However, discussing this aspect is beyond the scope of our study.
@@
Table 6. $imp_{VTS}$ and $imp_{VFS}$ for retrained and previous detectors.
@@
Figure 15. Examples of original fixed version and manually fixed version.
@@
We manually selected 100 TP samples and their corresponding fixed versions from each sub-dataset of SARD. Various methods were employed to manually correct the vulnerable samples. For example, Fig 15 showcases a piece of patched vulnerable code example, “-” denoted removed lines from vulnerable code, “+” denotes added lines in the patched version. Where lines 3-5 are originally fixed ways. lines 6-8 and 9 are two different manually fixed ways. During the manual correction process, we ensured that all the tokens incorporated into the code were already present in the original SARD dataset to avoid the out-of-vocabulary (OOV) problem of the embedding models. Subsequently, we evaluated both corrected code versions using vulnerability detectors and compared their respective accuracies. Following experiments in section 5.5.1, we opt not to carry out this experiment on the Fan dataset for the same reason.
@@
Furthermore, during this experiment, we incidentally observe a text pattern capable of inducing misclassifications for DeepWuKong, as illustrated in Fig 16. It displays a vulnerable code snippet and its corresponding fixed version. During preprocessing for DeepWuKong, variable names are transformed into symbolic names, such as “VAR” followed by an index. However, issues in the preprocessing pipeline have resulted in different symbolic names being assigned to identical variables in the vulnerable and fixed codes. Surprisingly, DeepWuKong can still accurately predict the symbolized samples. To investigate further, we manually adjusted the symbolization by changing the variable order. Remarkably, DeepWuKong misclassifies both the vulnerable and fixed code in this scenario. This observation indicates that vulnerability detectors rely more on recognizing textual patterns rather than truly comprehending underlying vulnerability patterns. Consequently, this limitation may hinder the ability of explanation approaches to identify vulnerability lines.
@@
6. Discussion
6.1. Limitation
First, we are not rigorously studying whether these explanation methods descriptively accurately explain the model’s output. In our study, we attempted to employ perturbation-based methods to compute the fidelity of explanation results and assess the significance of VTS and VFS in determining the model’s importance. However, subsequent observations revealed a paradoxical decrease in fidelity despite the retrained model showing higher FLC. Moreover, the model is indeed susceptible to perturbations that can alter the predicted labels, regardless of vulnerabilities. As a consequence, using perturbation-based assessments to determine the significance of certain lines of code for the model holds limited interpretive value. Therefore, we can only infer, through the comparison between retrained and previous models, that apart from perturbation-based techniques, other explanation methods can, to some extent, elucidate the model’s reasoning process. However, it’s important to note the inherent limitations of this evaluation approach.
@@ 6.1. Limitation
Third, Our research does not involve vulnerability detection methods based on large language models  (LLMs). The detectors we employed are primarily based on initial embeddings generated through methods like Word2Vec (word2vec, ), followed by input into GNN or RNN for binary classification. However, with the introduction of LLMs such as CodeBert (codebert, ), significant transformations have emerged in the methodology of code embeddings. These large language models often exhibit enhanced efficiency when performing code embeddings compared to traditional methods like Word2Vec. Fu M et al. (linevul, ) demonstrated that LLMs outperform IVDetect. However, we have identified limitations in current binary classification-based detectors, which primarily concentrate on distinctions between vulnerable and normal samples.
@@ 6.2. Implications
  • Firstly, acquiring a high-quality vulnerability dataset is not straightforward. We note that models retrained on a more refined sub-dataset exhibit a greater focus on the VFS compared to the previous models. This indicates a relatively reduced susceptibility to artifacts unrelated to vulnerabilities. However, acquiring such high-quality datasets in real-world scenarios poses considerable challenges, as highlighted by R Croft et al. (dataQuality, ), who discovered error rates in labeling vulnerabilities ranging from 20% to 71% within existing real-world datasets.
@@
Insights for future researches. Therefore, performing vulnerability detection based on binary classification may not be an optimal or efficient approach. Note that Jing Kai et al. (RepresentationLearning, ) found that while deep learning-based methods perform well in code classification and clone detection tasks, their performance in vulnerability detection is unsatisfactory. This indicates a lack of effective learning of vulnerability semantics by these detectors. We suggest that future researchers consider subdividing vulnerability detection into multiple sub-tasks rather than persisting with methods solely through binary classification tasks. For instance, they could follow the approach proposed by Y. Sun et al (gptscan, ), which initially employs large language models (LLMs) to identify variables and statements within the programs that correlate with vulnerabilities. Subsequently, they can use existing static analysis techniques to verify properties such as accessibility between these variables and statements, thereby identifying vulnerabilities. This approach allows vulnerability detection tools to leverage the latest advancements in deep learning for code comprehension, particularly the utilization of LLMs. Simultaneously, it mitigates the shortcomings in the reasoning abilities of deep learning models. Ultimately, this method optimizes the utilization of deep learning for enhanced and more efficient vulnerability detection.
@@ 7.1. Vulnerability Detection
To address the limitation, researchers have proposed various approaches. Li Z et al. (vuldeelocator, ) introduce the concept of granularity refinement and present BRNN-vdl, a model that can identify vulnerable tokens within input programs. Fu M and Tantithamthavorn C (linevul, ) employ a fine-tuned CodeBert (codebert, ) model to predict vulnerable functions and utilize its inner attention mechanism to identify critical sections. D Hin et al. (linevd, ) tackle vulnerability detection as a node-level classification problem, where they directly predict vulnerable statements. These approaches aim to enhance the precision and accuracy of identifying specific vulnerability-related statements within code. However, These approaches rely on detection models that require specific models to work, thus lacking generality.
@@ 7.2. Explanation Approaches
Surrogate methods (LIME, ; graphlime, ; pgmexplainer, ; relex, ) presuppose that the complexities in the neighboring regions of the input example are relatively low and can be effectively represented by a simpler surrogate model. They usually train interpretable models like random forests (RF, ), to approximate the predictions of the black-box model. While perturbation-based methods (gnnexplainer, ; pgexplainer, ; SubgraphX, ; Zorro, ; CasualScreen, ; GraphMask, ; IDLP, ; VulExplainer, ) investigate how perturbations in input features impact the model’s output score. By masking out trivial features and focusing on critical ones, these methods aim to identify the important features contributing to the models’ decisions. Gradient-based methods (GradInput, ; grad, ; gradcam++, ) typically compute the importance of input features by analyzing gradients or hidden feature maps. Decomposition methods (LRP-LSTM, ; gnnlrp, ) assess the significance of input features by decomposing the initial model predictions into multiple terms.
@@ 7.3. Explainable AI for SE
In the field of software engineering, defect prediction stands out as one of the primary domains where explainability methods find extensive application. For instance, several defect prediction tools (defect1_wattanakriengkrai2020predicting, ; defect2_pornprasit2021jitline, ; defect3_zheng2021just, ) employ LIME (LIME, ) to interpret the predictions of deep learning models. To address the limitations of LIME, which relies on random perturbations and may yield suboptimal results, Chanathip et al. (pornprasit2021pyexplainer, ) have introduced PyExplainer to explain defect prediction models. Experimental results have proved its superiority over LIME. Jack Humphrey et al. (humphreys2019explainable, ) have achieved defect prediction model interpretation by directly incorporating self-attention mechanisms. In other subfields, MRI Rabin et al. (rabin2021understanding, ) have developed a model-agnostic interpretability method, named Sivand, for method name prediction and variable name misuse prediction models. The key insight lies in the notion that by eliminating extraneous portions of input programs from the prediction process, one can gain a more profound understanding of the essential features involved in model inference. Rabin (rabin2022syntax, ) applied the syntax-aware variant Sivand-Perses to the same code intelligence tasks. They observed that Sivand-Perses is not only faster but also produces more compact sets of critical tokens in the reduced programs. Furthermore, they found that these critical tokens could be employed in generating adversarial examples for as many as 65% of the input programs.
To evaluate the effectiveness of these explainers in the context of vulnerability detection, T Ganz et al. (ACMExplainDiscovery, ) conducted a study using various explanation approaches for graph neural networks. They proposed criteria such as stability and descriptiveness to assess the explainers. The study revealed that all evaluated explanation methods exhibited deficiencies in at least two of the criteria. Hu Y et al. (hu2023interpreters, ) extensively evaluated six popular interpreters on four graph-based vulnerability detectors. They found that these interpreters performed poorly in terms of effectiveness, stability, and robustness. Specifically, instance-independent methods outperformed others in effectiveness, perturbation-based methods showed greater stability, and instance-independent approaches provided more consistent results for similar vulnerabilities in terms of robustness. Compared to the previous two works, our research uncovers additional concerns within DL-based detectors. Specifically, we identify that these detectors, despite achieving accurate detection performance, may still learn features unrelated to vulnerabilities. Moreover, they are susceptible to being misled by irrelevant perturbations, leading to a reversal in predictions, thereby influencing the effectiveness of explanation methods.

### 21 Feb 2024 v2

@@ 5.4.4. FLC Among Sequence-based Explainers
Fig. 10 illustrates the FLC results of sequence-based explainers, it reveals that:
The FLC scores achieved by sequence-based explainers fall short of expectations. On the Fan dataset, when explaining TokenLSTM, Shap attains the best FLC score, while it only reaches 30% of Expectation as shown in Fig. 10 (b.1). On the SARD dataset, Shap performs the best in FLC among the five explainers, however, the observed outcome diverges significantly from Expectation as depicted own in Fig. 10 (a.1),(a.2),(a.3).
There is minimal correspondence between the fidelity scores and FLC. One example is that Shap continues to outperform other explainers in terms of FLC on the SARD dataset. However, it does not achieve the highest fidelity scores across all detectors as shown in Fig. 7 (a.1),(b.1),(c.1).
Considering the aforementioned observations, it is uncertain whether explanation methods can reliably provide valuable insights. We hypothesize that DL-based detectors aim to learn the distinctions between vulnerable and normal samples, attempting to discern whether the content represents a vulnerability or normal code fragments. Given that VTSs exist in both vulnerable and non-vulnerable programs, DL-based detectors might not prioritize the importance of VTS in determining vulnerabilities. Hence, theoretically, explanation methods should achieve high FLC. As they focus on the differences between vulnerable and normal samples. However, experimental outcomes do not consistently align with this expectation. Furthermore, there appears to be no direct correlation observed among fidelity, TLC, and FLC values. As suggested by T Ganz (ACMExplainDiscovery, ), this discrepancy might stem from DL-based detectors being trained not to recognize vulnerabilities but artifacts encompassing not only VFS. These findings underscore the need for further analysis of DL-based detectors and explanation approaches, which will be discussed in § 5.5.
@@ 5.5.1.1   Experimental Setting
Dataset and detector consideration. When sampling the dataset, we chose to sample from the SARD portion and not from the Fan dataset. This decision was influenced by concerns regarding the potential presence of vulnerability-unrelated code changes in patch commits, as discussed in (dataQuality, ). Such changes could result in inaccurate labeling of vulnerabilities within the dataset. Simultaneously, we solely employ function-level detectors for experimentation and refrain from using slice-level detectors. This decision is based on the fact that DeepWuKong, SySeVR, and VulDeePecker detectors conduct slicing and symbolization operation. This process potentially exacerbates the differences unrelated to vulnerabilities between vulnerable and normal samples. For example, during symbolization, variable names such as “data” or “databuffer” in the original samples may be symbolized as “VAR1” and “VAR2”. However, in different samples or between a vulnerability sample and its corresponding normal sample, these variables may be symbolized as different values, such as “VAR1” in the vulnerability sample and “VAR2” in the normal sample. Subsequent experiments (§ 5.5.2) reveal that these textual disparities interfere to some extent with DL-based detectors.
@@
Dataset and detector selection. After sampling, the remaining vulnerable sample sizes in the sub-test set are shown in Table 3, indicating a considerable reduction in the dataset sizes for each type of vulnerability. The most notable reductions are observed in CWE-190 and CWE-400. On the sub-dataset, we first retrain the models for Reveal, IVDetect, Devign, and TokenLSTM. The comparison of the performance on the sub-test dataset between the previous and retrained models for each detector is depicted in Fig. 11. We can observe that Reveal and IVDetect exhibit minor performance differences between the previous and retrained models across the six types of sub-test datasets. However, Devign shows performance declines in CWE-20, CWE-119, and CWE-787, while TokenLSTM specifically demonstrates reduced performance in CWE-125, CWE-190, and CWE-400. We speculate that the performance declines may be related to the quantity and quality of the training samples. Considering the requirement for similar performance between the previous and retrained models, the number of test samples, and the need to involve both graph-based and sequence-based methods, we opt to continue the experiments on portions of CWE-20 and CWE-119 samples for Reveal, IVDetect, and TokenLSTM.
@@
Table 3. Number of vulnerable functions in each testset after filtering.
@@ 5.5.1.2   Experimental Results
The explanation results for Reveal, IVDetect, and TokenLSTM are depicted in Fig. 1213 and  14 respectively. From the results, we can observe:
@@
DL-based detectors are focused on differentiating vulnerable and normal samples, a characteristic also mirrored by gradient-, decomposition-, and surrogate-based explainers. Within Reveal, all model-related explainers demonstrate an ability to achieve higher FLC scores while concurrently registering decreased TLC. Regarding IVDetect, while the FLC score of DeepLift remains relatively modest without notable improvements, both GradCAM and GNN-LRP consistently maintain notably high FLC scores, exhibiting a slight increase, particularly evident in the retrained model. In the case of TokenLSTM, it is notable that all deployed explainers, with the exception of DeepLift, attain higher FLC scores in the retrained models. The aforementioned observations suggest that the retrained model indeed tends to place more emphasis on the VFS during predictions compared to the previous model. Additionally, explanation methods, beyond perturbation-based approaches, to some extent, can capture this particular aspect of the model. The results depicted in Table 6 reveal that in both the retrained and previous models, the significance of VFS surpasses that of VTS. Nevertheless, when examining their explanations in both retrained and previous models, their results may still not be considered ideal. This implies that they are still prone to selecting features unrelated to vulnerabilities, which could be associated with both the deep learning models and the explanation methods themselves. However, discussing this aspect is beyond the scope of our study.
@@
Table 6. $imp_{VTS}$ and $imp_{VFS}$ for retrained and previous detectors. Higher $imp$ scores reflect the significance of VTS/VFS in model decision-making.
@@
Figure 15. Examples of the original fixed version and manually fixed version.
@@
We manually selected 100 TP samples and their corresponding fixed versions from each sub-dataset of SARD. Various methods were employed to manually correct the vulnerable samples. For example, Fig. 15 showcases a piece of patched vulnerable code example, “-” denoted removed lines from vulnerable code, “+” denotes added lines in the patched version. Where lines 3-5 are originally fixed ways. lines 6-8 and 9 are two different manually fixed ways. During the manual correction process, we ensured that all the tokens incorporated into the code were already present in the original SARD dataset to avoid the out-of-vocabulary (OOV) problem of the embedding models. Subsequently, we evaluated both corrected code versions using vulnerability detectors and compared their respective accuracies. Following experiments in § 5.5.1, we opt not to carry out this experiment on the Fan dataset for the same reason.
@@
Furthermore, during this experiment, we incidentally observe a text pattern capable of inducing misclassifications for DeepWuKong, as illustrated in Fig. 16. It displays a vulnerable code snippet and its corresponding fixed version. During preprocessing for DeepWuKong, variable names are transformed into symbolic names, such as “VAR” followed by an index. However, issues in the preprocessing pipeline have resulted in different symbolic names being assigned to identical variables in the vulnerable and fixed codes. Surprisingly, DeepWuKong can still accurately predict the symbolized samples. To investigate further, we manually adjusted the symbolization by changing the variable order. Remarkably, DeepWuKong misclassifies both the vulnerable and fixed code in this scenario. This observation indicates that vulnerability detectors rely more on recognizing textual patterns rather than truly comprehending underlying vulnerability patterns. Consequently, this limitation may hinder the ability of explanation approaches to identify vulnerability lines.
@@
6. Discussions
6.1. Limitations
First, we are not rigorously studying whether these explanation methods descriptively and accurately explain the models’ output. In our study, we attempted to employ perturbation-based methods to compute the fidelity of explanation results and assess the significance of VTS and VFS in determining the model’s importance. However, subsequent observations revealed a paradoxical decrease in fidelity despite the retrained model showing higher FLC. Moreover, the model is indeed susceptible to perturbations that can alter the predicted labels, regardless of vulnerabilities. As a consequence, using perturbation-based assessments to determine the significance of certain lines of code for the model holds limited interpretive value. Therefore, we can only infer, through the comparison between retrained and previous models, that apart from perturbation-based techniques, other explanation methods can, to some extent, elucidate the model’s reasoning process. However, it’s important to note the inherent limitations of this evaluation approach.
@@ 6.1. Limitations
Third, Our research does not involve vulnerability detection methods based on large language models  (LLMs). The detectors we employed are primarily based on initial embeddings generated through methods like Word2Vec (word2vec, ), followed by input into GNN or RNN for binary classification. However, with the introduction of LLMs such as CodeBert (codebert, ), significant transformations have emerged in the methodology of code embeddings. These large language models often exhibit enhanced efficiency when performing code embeddings compared to traditional methods like Word2Vec. Fu et al. (linevul, ) demonstrated that LLMs outperform IVDetect. However, we have identified limitations in current binary classification-based detectors, which primarily concentrate on distinctions between vulnerable and normal samples.
@@ 6.2. Implications
  • Firstly, acquiring a high-quality vulnerability dataset is not straightforward. We note that models retrained on a more refined sub-dataset exhibit a greater focus on the VFS compared to the previous models. This indicates a relatively reduced susceptibility to artifacts unrelated to vulnerabilities. However, acquiring such high-quality datasets in real-world scenarios poses considerable challenges, as highlighted by Croft et al. (dataQuality, ), who discovered error rates in labeling vulnerabilities ranging from 20% to 71% within existing real-world datasets.
@@
Insights for future research. Therefore, performing vulnerability detection based on binary classification may not be an optimal or efficient approach. Note that Kai et al. (RepresentationLearning, ) found that while DL-based methods perform well in code classification and clone detection tasks, their performance in vulnerability detection is unsatisfactory. This indicates a lack of effective learning of vulnerability semantics by these detectors. We suggest that future work consider dividing vulnerability detection into multiple sub-tasks rather than persisting with methods solely through binary classification tasks. For instance, they could follow the approach proposed by Sun et al. (gptscan, ), which initially employs large language models (LLMs) to identify variables and statements within the programs that correlate with vulnerabilities. Subsequently, they can use existing static analysis techniques to verify properties such as accessibility between these variables and statements, thereby identifying vulnerabilities. This approach allows vulnerability detection tools to leverage the latest advancements in deep learning for code comprehension, particularly the utilization of LLMs. Simultaneously, it mitigates the shortcomings in the reasoning abilities of deep learning models. Ultimately, this method optimizes the utilization of deep learning for enhanced and more efficient vulnerability detection.
@@ 7.1. Vulnerability Detection
To address the limitation, researchers have proposed various approaches. Li et al. (vuldeelocator, ) introduce the concept of granularity refinement and present BRNN-vdl, a model that can identify vulnerable tokens within input programs. Fu et al. (linevul, ) employ a fine-tuned CodeBert (codebert, ) model to predict vulnerable functions and utilize its inner attention mechanism to identify critical sections. Hin et al. (linevd, ) tackle vulnerability detection as a node-level classification problem, where they directly predict vulnerable statements. These approaches aim to enhance the precision and accuracy of identifying specific vulnerability-related statements within code. However, these approaches rely on detection models that require specific models to work, thus lacking generality.
@@ 7.2. Explanation Approaches
Surrogate methods (LIME, ; graphlime, ; pgmexplainer, ; relex, ) assume that the complexities in the neighboring regions of the input example are relatively low and can be effectively represented by a simpler surrogate model. They usually train interpretable models like random forests (RF, ), to approximate the predictions of the black-box model. While perturbation-based methods (gnnexplainer, ; pgexplainer, ; SubgraphX, ; Zorro, ; CasualScreen, ; GraphMask, ; IDLP, ; VulExplainer, ) investigate how perturbations in input features impact the model’s output score. By masking out trivial features and focusing on critical ones, these methods aim to identify the important features contributing to the models’ decisions. Gradient-based methods (grad, ; gradcam++, ) typically compute the importance of input features by analyzing gradients or hidden feature maps. Decomposition methods (LRP-LSTM, ; gnnlrp, ) assess the significance of input features by decomposing the initial model predictions into multiple terms.
@@ 7.3. Explainable AI for SE
In the field of software engineering, defect prediction stands out as one of the primary domains where explainability methods find extensive application. For instance, several defect prediction tools (defect1_wattanakriengkrai2020predicting, ; defect2_pornprasit2021jitline, ; defect3_zheng2021just, ) employ LIME (LIME, ) to interpret the predictions of deep learning models. To address the limitations of LIME, which relies on random perturbations and may yield suboptimal results, Chanathip et al. (pornprasit2021pyexplainer, ) have introduced PyExplainer to explain defect prediction models. Experimental results have proved its superiority over LIME. Humphrey et al. (humphreys2019explainable, ) have achieved defect prediction model interpretation by directly incorporating self-attention mechanisms. In other subfields, Rabin et al. (rabin2021understanding, ) have developed a model-agnostic interpretability method, named Sivand, for method name prediction and variable name misuse prediction models. The key insight lies in the notion that by eliminating extraneous portions of input programs from the prediction process, one can gain a more profound understanding of the essential features involved in model inference. Rabin (rabin2022syntax, ) applied the syntax-aware variant Sivand-Perses to the same code intelligence tasks. They observed that Sivand-Perses is not only faster but also produces more compact sets of critical tokens in the reduced programs. Furthermore, they found that these critical tokens could be employed in generating adversarial examples for as many as 65% of the input programs.
To evaluate the effectiveness of these explainers in the context of vulnerability detection, Ganz et al. (ACMExplainDiscovery, ) conducted a study using various explanation approaches for graph neural networks. They proposed criteria such as stability and descriptiveness to assess the explainers. The study revealed that all evaluated explanation methods exhibited deficiencies in at least two of the criteria. Hu et al. (hu2023interpreters, ) extensively evaluated six popular interpreters on four graph-based vulnerability detectors. They found that these interpreters performed poorly in terms of effectiveness, stability, and robustness. Specifically, instance-independent methods outperformed others in effectiveness, perturbation-based methods showed greater stability, and instance-independent approaches provided more consistent results for similar vulnerabilities in terms of robustness. Compared to the previous two works, our research uncovers additional concerns within DL-based detectors. Specifically, we identify that these detectors, despite achieving accurate detection performance, may still learn features unrelated to vulnerabilities. Moreover, they are susceptible to being misled by irrelevant perturbations, leading to a reversal in predictions, thereby influencing the effectiveness of explanation methods.